# From minimal-length quantum theory to modified gravity

### v1

@@ I Introduction
The search for observable imprints of quantum gravity in semi-classical regimes has led to the development of various phenomenological models. Among them, the generalized uncertainty principle (GUP) has emerged as a prominent candidate for encoding Planck-scale corrections to quantum mechanics. By modifying the standard Heisenberg uncertainty relation, the GUP introduces a momentum-dependent deformation that incorporates a minimal measurable length [Kempf:1994su, Nozari:2005mr, Das:2009hs, Pedram:2011xj, Hossenfelder:2012jw, Bosso:2023sxr]. This minimal scale, often associated with the Planck length, is a common feature in several approaches to quantum gravity, including string theory, loop quantum gravity and non-commutative geometry [Gross:Mende:1988, AmelinoCamelia:2002dsr, Rovelli:1998lrr, Maggiore:1993rv, Konishi:1989wk, Capozziello:1999wx, Bosso:2024nmn].
On the gravitational side, classical general relativity (GR) can be generalized by including higher-order curvature invariants in the action. Such extensions give rise to theories such as $f(R)$ gravity [Starobinsky:1980te, Barrow:1988xh, Capozziello:2003tk, Clifton:2005aj, Starobinsky:2007hu, Sotiriou:2008rp, DeFelice:2010aj, Nojiri:2017ncd, Capozziello:2019cav, Bajardi:2022ocw], or more broadly to Lagrangians depending on both the Ricci scalar and tensor invariants, as in $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity [Carroll:2004de, Bogdanos:2009tn, Bajardi:2023shq]. These modifications are not introduced arbitrarily; rather, they emerge naturally within effective field theory treatments of gravity coupled to quantum matter fields [Birrell:1982ix], and they are supported by the renormalization structure of quantum field theory in curved spacetime [Sotiriou:2008rp, Nojiri:2010wj].
@@
Several studies have investigated how GUP-induced corrections affect black hole (BH) physics [Scardigli:1999jh, Buoninfante:2019fwr, Adler:2001vs, Medved:2004yu, Custodio:2003jp], with particular focus on the entropy–area relation [Majumder:2011wl, Bosso:2023fnb, Buoninfante:2020guu, Tawfik:2015kga]. Such corrections typically involve logarithmic and inverse-area terms, which are likewise predicted by higher-curvature gravity theories. This formal resemblance suggests a deeper structural connection between the two frameworks and motivates a more systematic investigation of their correspondence. Despite these advances, a general methodology for reconstructing effective gravitational Lagrangians from arbitrary GUP deformations remains lacking. Most existing analyses, in fact, focus on particular GUP forms and a limited class of gravitational models [Hammad:2015dka], often without providing a unified procedure that connects the deformation function to curvature corrections at the action level.
In this paper, we propose a general framework to establish this connection. Starting from a generic expansion of the GUP deformation function in powers of momentum uncertainty, we derive the corresponding corrections to BH entropy by interpreting the GUP-modified uncertainty as a regulator near the horizon. These entropy corrections are then translated into modifications of the gravitational action using Wald’s entropy formula, applied first to $f(R)$ and then to the more general $f(R,R_{\mu\nu}R^{\mu\nu})$ class of theories. By inverting the resulting relations, we obtain an explicit reconstruction of the effective Lagrangian, identifying the curvature-dependent terms that encode the underlying quantum uncertainty structure. As such, our approach provides a clear and physically motivated dictionary connecting quantum mechanical deformations to semi-classical gravitational dynamics. Furthermore, it represents a natural and direct extension of the results of [Hammad:2015dka] to the fully general case of polynomial GUP models with deformation functions of arbitrary degree.
@@
Throughout this work, we adopt natural units, $k_{B}=\hbar=c=1$, such that the Planck length is expressed as $\ell_{P}=\sqrt{G}$, where $G$ denotes Newton’s gravitational constant.
@@ II Generalized Uncertainty Principle and Black hole Entropy
The GUP is a phenomenological modification of the standard Heisenberg uncertainty relation that incorporates effects expected from quantum gravity theories. It generally introduces a minimal measurable length scale, typically on the order of the Planck length, which modifies the standard commutation relations between position and momentum operators [Kempf:1994su].
A commonly studied class of models introduces deformations that depend on powers of the momentum uncertainty or, alternatively, on the momentum operator, with the two formulations related through the Robertson-Schrödinger relation. A well-known example is the model proposed in [Kempf:1994su], in which the canonical commutator is deformed as $x,p]=i(1+\beta p^{2})$, where $\beta$ is a small deformation parameter, typically associated with the (squared) Planck scale. This modification leads to an uncertainty relation of the form $\Delta x\Delta p\geq\frac{1}{2}\left[1+\beta(\Delta p)^{2}\right]$, which implies the existence of a non-zero minimal uncertainty in position, $\Delta x_{\text{min}}=\sqrt{\beta}>0$. The physical consequences of this feature have been widely discussed in the literature [Scardigli:2003kr, Jizba:2009qf, Amelino-Camelia:2005zpp, Das:2008kaa, Carr:2015nqa, Scardigli:2018jlm, Anacleto:2015mma, Jizba:2022icu, Tawfik:2015rva, Barca:2023shv, Segreto:2024gcg] (see also [Bosso:2023aht] for a recent review).
Alternative formulations based on deformed commutators have also been proposed, such as $x,p]=i\,e^{\beta p^{2}} [$Nouicer:2007jg] or $x,p]=i(1-\beta p^{2})^{-1} [$Pedram:2011gw], which represent non-polynomial generalizations of the canonical commutation relation and lead to qualitatively different physical implications. The use of even powers, such as $p^{2}$ or $(\Delta p)^{2}$, is motivated primarily by symmetry considerations. In particular, these functions are invariant under parity transformations $p\rightarrow-p$, preserving the isotropy of space and ensuring that the deformation respects fundamental discrete symmetries. However, more exotic models have been put forward in which the deformation also includes odd powers of the momentum, such as a linear dependence [Ali:2011fa]. These terms, which explicitly break parity symmetry, may emerge in theories with intrinsic chirality, anisotropic backgrounds or non-commutative geometries.
@@
In order to proceed with analytical calculations, it is customary to assume that inequality (1) is approximately saturated. This approximation corresponds to the coherent states of the ordinary theory, i.e., states characterized by a minimal uncertainty product. Such states, in the framework of minimal-length quantum mechanics, have been extensively considered in the literature (see, e.g., [Kempf:1994su, Pedram:2012ui, Dey:2012tv, Bosso:2017ndq, Bosso:2021koi, Jizba:2023ygi]). From a computational perspective, the use of an approximate equality is essential to derive closed-form expressions or to make progress in perturbative expansions. Moreover, many physical applications of the GUP, such as in BH thermodynamics, quantum cosmology or minimal length models, rely on this approximation to extract meaningful phenomenological consequences [Hossenfelder:2012jw].
@@ II.1 Implications for BH entropy
This constraint directly affects the properties of the BH horizon area. When a particle is absorbed, the minimal increase in the horizon area $\Delta A$ is linked to the particle’s position uncertainty $\Delta x$ [Awad:2014bta, Hammad:2015dka]. Incorporating the GUP-modified uncertainty leads to corrections in the quantization of horizon area and consequently to corrections in the Bekenstein-Hawking entropy formula:
@@
where $S_{\mathrm{corr}}$ contains sub-leading terms due to quantum gravity effects. On the other hand, starting from BH absorption and emission, with a Gedankenexperiment similar to Heisenberg’s microscope, one can arrive at a GUP formulation [Maggiore:1993rv].
As stated above, the heuristic derivation of $S_{\mathrm{corr}}$ proceeds by estimating the minimal area increase $\Delta A$ caused by absorption of a quantum particle with position uncertainty $\Delta x$ and momentum uncertainty $\Delta p$. Since the energy of a particle with size $\Delta x$ near the horizon can be estimated by $E\sim\Delta p$ and considering that the minimal increase in BH area associated with the absorption of such a particle is $\Delta A_{\text{min}}\sim GE\Delta x$ [Hammad:2015dka], we obtain from Eq. (4)
@@
Replacing $\Delta x=\zeta\sqrt{A}$, where $\zeta$ is a dimensionless proportionality constant, Eq. (8) can be expressed as a function of the horizon area:
@@
Nevertheless, if for each $k$ one considers only the leading corrections [^1] For example, if only $c_{1}\neq 0$, the leading-order term is given by Eq. (6a), as $c_{1}^{n}$ with $n>1$ are subdominant and, thus, $\tilde{c}_{k}=0,\,\forall k>1$., one may infer a direct correspondence between the terms in the GUP expansion and the corrections to the entropy. For instance, we observe that if $c_{k}=0,\,\forall k\neq 1$, the Bekenstein-Hawking entropy acquires a correction that scales as $\sqrt{A}$, which is indicative of a genuinely infrared (IR) effect. This outcome is consistent with results obtained in other approaches involving linear GUP deformations; see, for example, [Majumder:2012rtc, Tawfik:2015kga]. Furthermore, it has been identified in non-commutative quantization contexts [Nozari:2008rc]. It is therefore evident that, when focusing on the ultraviolet (UV) regime of quantum gravity, such a term becomes subdominant compared to the leading corrections and can be safely neglected. From the GUP perspective, this provides a further argument in support of models in which the linear term is absent and the leading-order deformation is quadratic in $\Delta p$ [Kempf:1994su], as these models are more naturally associated with the existence of a minimal length and are better suited to capturing the dominant features of the UV regime.
On the other hand, if only $c_{2}\neq 0$, we have a logarithmic modification of the entropy-area relation. This term emerges as a nearly universal feature of UV quantum gravity, although different approaches yield variations in the numerical coefficient that multiplies the logarithmic contribution [Carlip:2000nv, Kaul:2000kf, Sen:2012dw, Adler:2001vs, Medved:2004yu]. Additionally, this type of correction is supported by entanglement entropy calculations in BH spacetimes in the UV regime [Solodukhin:2011gn, Alonso-Serrano:2022qvo]. Finally, for each $c_{k}\neq 0$ with $k\geq 3$, the entropy corrections scale as $A^{1-k/2}$, indicating that they become relevant only in the deep UV regime.
@@ III Modified Gravity and Wald Entropy
represent a natural extension of GR. Here, $g$ is the determinant of the metric tensor, and the Lagrangian density $f(R)$ is a non-linear function of the Ricci scalar $R$, encoding higher-order curvature corrections expected from quantum gravitational or effective field theory considerations [Nojiri:2010wj, Sotiriou:2008rp].
Such extensions emerge naturally within various theoretical frameworks. For instance, quantum field theory formulated on curved spacetime induces one-loop quantum corrections that generate terms quadratic and of higher order in curvature invariants, as extensively discussed in [Birrell:1982ix]. Similarly, effective actions derived from string theory include $\alpha$-corrections, which are systematically expressed as powers of curvature tensors and represent higher-derivative modifications to classical gravity [Zwiebach:1985uq]. Additionally, phenomenological approaches aiming to explain cosmological phenomena such as dark energy and inflation often invoke modified gravitational actions featuring curvature-dependent terms [DeFelice:2010aj]. These higher-derivative corrections modify the gravitational dynamics at small (high curvature) scales and may encode signatures of the underlying quantum theory of gravity.
The Bekenstein-Hawking entropy formula is derived within the framework of classical GR. In generalized gravity theories, this relation does not hold straightforwardly due to additional curvature-dependent terms in the action. Wald formulated a general expression for BH entropy applicable to any diffeomorphism-invariant theory of gravity [Wald:1993nt]. For the class of $f(R)$ theories, the Wald entropy is given by [Iyer:1994ys]
@@
In spherically symmetric spacetimes, the Ricci scalar evaluated at the BH horizon $R_{H}$ typically scales inversely with the square of the horizon radius, $r_{H}$:
@@
Thus, by focusing on constant curvature solutions, the Wald entropy in $f(R)$ gravity can be expressed as [DAgostino:2024sgm]
@@
This expression makes explicit how GUP-induced entropy corrections translate into non-trivial curvature terms in the gravitational action. Thus, the quantum mechanical deformation encoded by $\Xi(\Delta p)$ corresponds semi-classically to a modified gravitational Lagrangian. It is worth reminding that the consistency of this picture requires the positivity of $f_{R}(R)$ to ensure a well-defined effective gravitational coupling and avoid ghost instabilities [Sotiriou:2008rp].

### v2

@@ I Introduction
The search for observable imprints of quantum gravity in semi-classical regimes has led to the development of various phenomenological models. Among them, the generalized uncertainty principle (GUP) has emerged as a prominent candidate for encoding Planck-scale corrections to quantum mechanics. By modifying the standard Heisenberg uncertainty relation, the GUP introduces a momentum-dependent deformation that incorporates a minimal measurable length Kempf et al. (1995); Nozari and Azizi (2006); Das and Vagenas (2009); Pedram et al. (2011); Hossenfelder (2013); Bosso et al. (2023b). This minimal scale, often associated with the Planck length, is a common feature in several approaches to quantum gravity, including string theory, loop quantum gravity and non-commutative geometry Gross and Mende (1988); Amelino–Camelia (2002); Rovelli (1998); Maggiore (1993); Konishi et al. (1990); Capozziello et al. (2000); Bosso (2024).
On the gravitational side, classical general relativity (GR) can be generalized by including higher-order curvature invariants in the action. Such extensions give rise to theories such as $f(R)$ gravity Starobinsky (1980); Barrow and Cotsakis (1988); Capozziello et al. (2003); Clifton and Barrow (2005); Starobinsky (2007); Sotiriou and Faraoni (2010); De Felice and Tsujikawa (2010); Nojiri et al. (2017); Capozziello et al. (2019); D’Agostino and Nunes (2020); Bajardi et al. (2022), or more broadly to Lagrangians depending on both the Ricci scalar and tensor invariants, as in $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity Carroll et al. (2005); Bogdanos et al. (2010); Bajardi and D’Agostino (2024). These modifications are not introduced arbitrarily; rather, they emerge naturally within effective field theory treatments of gravity coupled to quantum matter fields Birrell and Davies (1982), and they are supported by the renormalization structure of quantum field theory in curved spacetime Sotiriou and Faraoni (2010); Nojiri and Odintsov (2011).
@@
Several studies have investigated how GUP-induced corrections affect black hole (BH) physics Scardigli (1999); Buoninfante et al. (2019); Adler et al. (2001); Medved (2005); Custodio and Horvath (2003); Ong (2025), with particular focus on the entropy-area relation Majumder (2011); D’Agostino (2019); Bosso et al. (2024); Buoninfante et al. (2022); Tawfik and El Dahab (2015); Chagoya et al. (2025). Such corrections typically involve logarithmic and inverse-area terms, which are also predicted by higher-curvature gravity theories. This formal resemblance suggests a deeper structural connection between the two frameworks and motivates a more systematic investigation of their correspondence.
Despite these advances, a general methodology for reconstructing effective gravitational Lagrangians from arbitrary GUP deformations is still lacking. Most existing analyses focus on particular GUP forms and a limited class of gravitational models Hammad (2015); Wojnar (2024); Ali and Wojnar (2024), often without providing a unified procedure that connects the deformation function to curvature corrections at the action level. Some progress in this direction has been achieved, e.g., in Ref. Dias et al. (2016), but it is restricted to a linear approximation of the high-order gravity action and does not yield a fully general covariant formulation.
In this paper, we propose a general framework to establish a systematic connection between GUP-inspired quantum corrections and extended gravitational actions incorporating higher-curvature terms. Starting from a generic expansion of the GUP deformation function in powers of momentum uncertainty, we derive the corresponding corrections to BH entropy by interpreting the GUP-modified uncertainty as a regulator near the horizon. These entropy corrections are then translated into modifications of the gravitational action using Wald’s entropy formula, applied first to $f(R)$ and then to the more general $f(R,R_{\mu\nu}R^{\mu\nu})$ class of theories. By inverting the resulting relations, we obtain an explicit reconstruction of the effective Lagrangian, identifying the curvature-dependent terms that encode the underlying quantum uncertainty structure. As such, our approach provides a clear and physically motivated dictionary connecting quantum mechanical deformations to semi-classical gravitational dynamics. Furthermore, it represents a natural and direct extension of the results of Ref. Hammad (2015) to the fully general case of polynomial GUP models with deformation functions of arbitrary degree.
@@
Throughout this work, we adopt natural units, $k_{B}=\hbar=c=1$, such that the Planck length is $\ell_{P}=\sqrt{G}$, where $G$ denotes Newton’s gravitational constant.
@@ II Generalized Uncertainty Principle and Black hole Entropy
The GUP is a phenomenological modification of the standard Heisenberg uncertainty relation that incorporates effects expected from quantum gravity theories. It generally introduces a minimal measurable length scale, typically on the order of the Planck length, which modifies the standard commutation relations between position and momentum operators Kempf et al. (1995).
A commonly studied class of models introduces deformations that depend on powers of the momentum uncertainty or, alternatively, on the momentum operator, with the two formulations related through the Robertson-Schrödinger relation. A well-known example is the model proposed in Ref. Kempf et al. (1995), in which the canonical commutator is deformed as $x,p]=i(1+\beta p^{2})$, where $\beta$ is a small deformation parameter, typically associated with the (squared) Planck scale. This modification leads to an uncertainty relation of the form $\Delta x\Delta p\geq\frac{1}{2}\left[1+\beta(\Delta p)^{2}\right]$, which implies the existence of a non-zero minimal uncertainty in position, $\Delta x_{\text{min}}=\sqrt{\beta}>0$. The physical consequences of this feature have been widely discussed in the literature Scardigli and Casadio (2003); Jizba et al. (2010); Amelino-Camelia et al. (2006); Das and Vagenas (2008); Carr et al. (2015); Scardigli et al. (2018); Anacleto et al. (2015); Jizba et al. (2022); Tawfik and Diab (2015); Barca and Montani (2024); Segreto and Montani (2025) (see also Bosso et al. (2023a) for a recent review).
Alternative formulations based on deformed commutators have also been proposed, such as $x,p]=i\,e^{\beta p^{2}}$ Nouicer (2007) or $x,p]=i(1-\beta p^{2})^{-1}$ Pedram (2012), which represent non-polynomial generalizations of the canonical commutation relation and lead to qualitatively different physical implications. The use of even powers, such as $p^{2}$ or $(\Delta p)^{2}$, is motivated primarily by symmetry considerations. In particular, these functions are invariant under parity transformations $p\rightarrow-p$, preserving the isotropy of space and ensuring that the deformation respects fundamental discrete symmetries. However, more exotic models have been put forward in which the deformation also includes odd powers of the momentum, such as a linear dependence Ali et al. (2011). These terms, which explicitly break parity symmetry, may emerge in theories with intrinsic chirality, anisotropic backgrounds or non-commutative geometries.
@@
In order to proceed with analytical calculations, it is customary to assume that inequality (1) is approximately saturated. This approximation corresponds to the coherent states of the ordinary theory, i.e., states characterized by a minimal uncertainty product. Such states, in the framework of minimal-length quantum mechanics, have been extensively considered in the literature (see, e.g., Refs. Kempf et al. (1995); Pedram (2013); Dey and Fring (2012); Bosso et al. (2017); Bosso and Luciano (2021); Jizba et al. (2023)). From a computational perspective, the use of an approximate equality is essential to derive closed-form expressions or to make progress in perturbative expansions. Moreover, many physical applications of the GUP, such as in BH thermodynamics, quantum cosmology or minimal length models, rely on this approximation to extract meaningful phenomenological consequences Hossenfelder (2013).
@@ II.1 Implications for BH entropy
This constraint directly affects the properties of the BH horizon area. When a particle is absorbed, the minimal increase in the horizon area $\Delta A$ is linked to the particle’s position uncertainty $\Delta x$ Awad and Ali (2014); Hammad (2015). Incorporating the GUP-modified uncertainty leads to corrections in the quantization of horizon area and consequently to corrections in the Bekenstein-Hawking entropy formula:
@@
where $S_{\mathrm{corr}}$ contains sub-leading terms due to quantum gravity effects. On the other hand, starting from BH absorption and emission, with a Gedankenexperiment similar to Heisenberg’s microscope, one can arrive at a GUP formulation Maggiore (1993).
As stated above, the heuristic derivation of $S_{\mathrm{corr}}$ proceeds by estimating the minimal area increase $\Delta A$ caused by absorption of a quantum particle with position uncertainty $\Delta x$ and momentum uncertainty $\Delta p$. Since the energy of a particle with size $\Delta x$ near the horizon can be estimated by $E\sim\Delta p$ and considering that the minimal increase in BH area associated with the absorption of such a particle is $\Delta A_{\text{min}}\sim GE\Delta x$ Hammad (2015), we obtain from Eq. (4)
@@
We remark that the relation $\Delta A_{\text{min}}\sim GE\Delta x$ should be understood as a rather general semiclassical, near-horizon estimate. Indeed, the area-radius relation $A=4\pi r_{H}^{2}$ is purely geometric and theory-independent; on the other hand, the absorption of a quantum with conserved energy $E$ induces a corresponding variation of the BH mass and horizon radius. In extended/modified gravity theories, this response may differ from the Einstein case through an effective gravitational coupling, while remaining linear in the absorbed energy. Consequently, one generically expects $\Delta A_{\text{min}}\propto G_{\mathrm{eff}}E\Delta x$, with $\Delta x\sim\mathcal{O}(r_{H})$ and the proportionality factor depending on the underlying gravitational dynamics Jacobson et al. (1995); Padmanabhan (2010). In our case, we absorb this factor into the coefficient $\lambda$.
Let us now adopt the usual near-horizon identification $\Delta x=\zeta\sqrt{A}$ Hammad (2015), where $\zeta$ is a dimensionless proportionality constant. This assumption is appropriate for uncharged and static BHs, where $\sqrt{A}\propto r_{H}$ is the only characteristic length scale of the system. Nevertheless, in the presence of additional degrees of freedom (e.g., charge, rotation, or higher-curvature couplings), one may consider the more general form $\Delta x=r_{H}\mathfrak{F}$, where $\mathfrak{F}$ is a dimensionless function depending on the characteristic parameters of the BH. Such dependence is expected to be reabsorbed through a redefinition of $\zeta$ without affecting the functional structure of the GUP-induced corrections. Under our assumption, Eq. (8) can be written as a function of the horizon area:
@@
Nevertheless, if for each $k$ one considers only the leading corrections [^1] For example, if only $c_{1}\neq 0$, the leading-order term is given by Eq. (6a), as $c_{1}^{n}$ with $n>1$ are subdominant and, thus, $\tilde{c}_{k}=0,\,\forall k>1$., one may infer a direct correspondence between the terms in the GUP expansion and the corrections to the entropy. For instance, we observe that if $c_{k}=0,\,\forall k\neq 1$, the Bekenstein-Hawking entropy acquires a correction that scales as $\sqrt{A}$, which is indicative of a genuinely infrared (IR) effect. This outcome is consistent with results obtained in other approaches involving linear GUP deformations Majumder (2013); Tawfik and El Dahab (2015). Furthermore, it has been identified in non-commutative quantization contexts Nozari and Mehdipour (2008). It is therefore evident that, when focusing on the ultraviolet (UV) regime of quantum gravity, such a term becomes subdominant compared to the leading corrections and can be safely neglected. From the GUP perspective, this provides a further argument in support of models in which the linear term is absent and the leading-order deformation is quadratic in $\Delta p$ Kempf et al. (1995), as these models are more naturally associated with the existence of a minimal length and are better suited to capturing the dominant features of the UV regime.
On the other hand, if only $c_{2}\neq 0$, we have a logarithmic modification of the entropy-area relation. This term emerges as a nearly universal feature of UV quantum gravity, although different approaches yield variations in the numerical coefficient that multiplies the logarithmic contribution Carlip (2000); Kaul and Majumdar (2000); Sen (2013); Adler et al. (2001); Medved (2005). Additionally, this type of correction is supported by entanglement entropy calculations in BH spacetimes in the UV regime Solodukhin (2011); Alonso-Serrano and Liška (2023). Finally, for each $c_{k}\neq 0$ with $k\geq 3$, the entropy corrections scale as $A^{1-k/2}$, indicating that they become relevant only in the deep UV regime.
@@ III Modified Gravity and Wald Entropy
represent a natural extension of GR. Here, $g$ is the determinant of the metric tensor, and the Lagrangian density $f(R)$ is a non-linear function of the Ricci scalar $R$, encoding higher-order curvature corrections expected from quantum gravitational or effective field theory considerations Nojiri and Odintsov (2011); Sotiriou and Faraoni (2010).
Such extensions emerge naturally within various theoretical frameworks. For instance, quantum field theory formulated on curved spacetime induces one-loop quantum corrections that generate terms quadratic and of higher order in curvature invariants, as extensively discussed in Ref. Birrell and Davies (1982). Similarly, effective actions derived from string theory include $\alpha$-corrections, which are systematically expressed as powers of curvature tensors and represent higher-derivative modifications to classical gravity Zwiebach (1985). Additionally, phenomenological approaches aiming to explain cosmological phenomena such as dark energy and inflation often invoke modified gravitational actions featuring curvature-dependent terms De Felice and Tsujikawa (2010). These higher-derivative corrections modify the gravitational dynamics at small (high curvature) scales and may encode signatures of the underlying quantum theory of gravity.
The Bekenstein-Hawking entropy formula is derived within the framework of classical GR. In generalized gravity theories, this relation does not hold straightforwardly due to additional curvature-dependent terms in the action. Wald formulated a general expression for BH entropy applicable to any diffeomorphism-invariant theory of gravity Wald (1993). For the class of $f(R)$ theories, the Wald entropy is given by Iyer and Wald (1994)
@@
In modified gravity theories, such as $f(R)$ models, where static BH solutions may exhibit a non-vanishing Ricci scalar Multamaki and Vilja (2006); de la Cruz-Dombriz et al. (2009), it is convenient to parametrize the horizon curvature in terms of the geometric scale given by the horizon radius. Following Ref. Hammad (2015), we thus assume the scaling relation [^2] Equation (18) should not be intended as a universal relation for the Ricci scalar valid for any BH geometries, but rather as a phenomenological ansatz appropriate for effective curvature corrections emerging in extended gravity theories.
@@
Thus, by focusing on constant curvature solutions, the Wald entropy in $f(R)$ gravity can be expressed as D’Agostino and Luciano (2024)
@@
This expression makes explicit how GUP-induced entropy corrections translate into non-trivial curvature terms in the gravitational action. Thus, the quantum mechanical deformation encoded by $\Xi(\Delta p)$ corresponds semi-classically to a modified gravitational Lagrangian. It is worth reminding that the consistency of this picture requires the positivity of $f_{R}(R)$ to ensure a well-defined effective gravitational coupling and avoid ghost instabilities Sotiriou and Faraoni (2010).